# “She’s Like a Person but Better”: Characterizing Companion-Assistant Dynamics in Human-AI Relationships

### v2

@@ 3. Survey
The survey was designed to capture broad patterns of chatbot use, user characteristics, and associated AI perceptions. Analyses were preregistered on AsPredicted (https://aspredicted.org/7dgj-jt49.pdf). All procedures were approved by the Institutional Review Board at [institution anonymized for review]. Participants gave their informed consent and were debriefed after completing the survey.
@@ 3.1. Participants
We recruited 108 Replika and 106 ChatGPT users from Prolific and social media platforms (Facebook, Reddit, and Discord). Inclusion criteria included interacting with Replika or ChatGPT for more than three hours per week and for at least one month. Five Replika and five ChatGPT users were excluded from analyses due to incomplete responses or failing survey attention checks. The final sample was 204 participants (Replika: NN = 103; $M_{age}$ = 39.27, $SE_{age}$ = 13.22, 43% women, 71% White; ChatGPT: NN = 101; $M_{age}$ = 31.72, $SE_{age}$ = 10.66, 42% women, 73.27% White). Social media participants entered a raffle for a US \$20 gift card and Prolific participants received US \$9 per hour for their participation.
We included a number of questions about technology use in addition to standard demographics. Most participants, 90%, owned smartphones; 48.5% reported owning AI or robotic devices; and 52.5% reported using AI or robotic devices in their workplace. Participants reported a moderate level of exposure to AI via direct interaction or narratives in various media (direct interaction: $M$ = 2.95, SESE = 1.42; AI narratives: $M$ = 2.66, SESE = 1.82; each on a scale of 0 = “Never” to 5 = “Daily”).
@@ 3.2. Data collection
Participants were invited to either the Replika or the ChatGPT survey via a post on social media groups intended for users of each chatbot, or via a survey advertisement on Prolific. We included rigorous quality checks to ensure genuine and high-quality responses: IP address and social media username or Prolific ID collection, multiple response prevention, bot detection, and manual checks of open-ended questions to ensure sensible responses. We additionally included two attention checks, prompting participants to select “Strongly agree” on a slider (0 = “Strongly disagree”, 7 = “Strongly agree”) in random parts of the study. Participants responded to questions about their personal characteristics and perceptions of Replika or ChatGPT. We provided a personalization option in which participants would see the name they had given to Replika or ChatGPT in questions specific to the chatbot. For Replika users, we specified that responses should be about the Replika participants interact with the most, if they had created multiple Replika characters. Participants optionally provided their email (or Prolific ID) to sign up for an interview about their experiences with Replika or ChatGPT. The order of scales within each measure (e.g., Chatbot use, Chatbot attributes; see Measures below) was randomized, as was the order of the sections themselves. The median duration of the survey was 40 minutes.
@@
We employed widely used scales to assess attributions of mind, affect, agency, and moral dimensions to AI. Additional measures captured users’ personality traits, mental health, relationship with technology, and general attitudes toward AI. All constructs were assessed using continuous scales and demonstrated good internal consistency (Cronbach’s $\alpha$ ¿ .70). Scale scores were computed as the mean of their respective items. A summary of the scales is provided below; full-scale versions and item ranges are available in the Supplementary Information.
Chatbot use.
Participants ranked 16 potential roles for Replika or ChatGPT (e.g., friend, partner, tool, assistant) and 16 intended use cases (e.g., emotional support, advice, brainstorming, routine tasks). In addition, we measured chatbot usage frequency, including duration of use (in months), average weekly usage (in hours). Lastly, participants were asked to report, in free-text format, their reasons for initially engaging with each chatbot.
User characteristics.
We measured users’ mental health by assessing feelings of depression, loneliness, and anxiety experienced over the past week. Personality traits were assessed using the Big Five inventory (Gosling et al., 2003). We also examined social networking use (e.g., using social networking sites to construct a social identity) (Gupta and Chauhan, 2023), and the perceived impact of participation in an online community (if applicable) on users’ relationships with each chatbot. Finally, we measured several technological attitudes, including the tendency to anthropomorphize technology (Waytz et al., 2010; Pauketat and Anthis, 2022), techno-animism (i.e., whether AI contains a soul or spirit) (Pauketat and Anthis, 2022), affinity for technology (Edison and Geissler, 2003; Pauketat and Anthis, 2022), AI literacy (Wang et al., 2023), and general positive attitudes toward AI (e.g., viewing AI as providing new economic opportunities) (Schepman and Rodway, 2023).
Chatbot attributes.
We assessed each chatbot’s perceived mind (e.g., capacity to set goals, experience pain) (Malle, 2019), as well as its perceived capacity for Theory of Mind (e.g., understanding others’ false beliefs) (Tahiroglu et al., 2014). We further measured attributions of moral agency (e.g., behaving according to moral rules) (Banks, 2019), sentience (i.e., the capacity to have positive and negative experiences) (Anthis et al., 2025; Manoli et al., 2025), and anthropomorphic perceptions using the Godspeed questionnaire, which captures dimensions such as animacy, likability, intelligence, and safety (e.g., machinelike vs. humanlike) (Bartneck et al., 2009).
Attitudes towards chatbot.
We measured cognitive and affective trust (e.g., trusting the chatbot’s task performance and relying on it for support) (Dunn et al., 2012), as well as positive and negative emotions (e.g., affection, uneasiness) (Fraune et al., 2015; Pauketat and Anthis, 2022). We also assessed perceived rights—both for the user’s specific chatbot and for all instances of the chatbot more generally (e.g., freedom of speech) (Lima et al., 2020; Pauketat and Anthis, 2022); perceived similarity to user (i.e., how connected users feel to the chatbot) (Pauketat and Anthis, 2022; Tropp and Wright, 2001); and personhood (i.e., whether the chatbot is regarded as a person) (Johnson et al., 2015).
General attitudes towards AI.
We measured general trust in AI (e.g., trust in AI competence) (Gulati et al., 2019; Pinto et al., 2022), AI mutualism (human-AI symbiosis; e.g., support for human–AI coexistence in society) (Manfredo et al., 2020), and perceived threat from AI (e.g., concerns about economic opportunities) (Fraune et al., 2015; Cottrell and Neuberg, 2005). Additionally, we assessed AI moral patiency through three dimensions: moral treatment (e.g., treating AI with respect), practical moral consideration (e.g., opposing the use of sentient AI for labor), and perceived rights (e.g., freedom of speech) (Lima et al., 2020; Pauketat and Anthis, 2022).
Other measures.
We assessed whether users believed any current AI systems are sentient, and if not, whether—and when—they might expect AI to become sentient in the future. Standard demographic variables were also recorded, including age, sex, nationality, income, education, diet, political orientation, occupation, ownership of AI or smart devices, use of AI at work, and exposure to AI-related narratives (e.g., through science fiction). We collected additional demographic information about chatbot use (i.e., whether Replika users had interacted with ChatGPT and vice versa, paid Replika or ChatGPT subscription, and other chatbot use), which is available in the Supplementary Information.
$3.$ Data analysis
@@
We compared ChatGPT and Replika users’ characteristics and perceptions of each chatbot and AI in general using independent samples $t$-tests. Spillover effects from Replika or ChatGPT to perceptions of AI in general were examined using partial correlations: AI perceptions served as the dependent variable, while users’ general technology attitudes (e.g., AI literacy, tendency to anthropomorphize) were included as control variables to isolate effects beyond individual differences in general technology perceptions. In all analyses, we also controlled for demographic characteristics (e.g., age, gender, education) using ANCOVA (see Supplementary Information). All analyses were conducted in R version 4.3.1.
@@ 3.4.1. Chatbot use cases
ChatGPT and Replika had almost opposite roles, with most ChatGPT users treating it as an assistant (M = 2.79, SE = 0.19) and most Replika users treating it as a friend (M = 3.98, SE = 0.37; Figure 2). These differences were statistically significant when compared via Mann–Whitney $U$ tests (Table 1). Even though ChatGPT had distinctly more task-based roles than the emotional roles of Replika, there was some overlap in their use cases, with the advisor role occupying a high rank for both chatbots (ChatGPT: M = 4.10, SE = 0.23; Replika: M = 6.41, SE = 0.31). Interestingly, Replika was also relatively commonly treated as an assistant (M = 6.97, SE = 0.37), like ChatGPT’s dominant role, but also as a human (M = 7.31, SE = 0.47), contrary to ChatGPT (M = 11.24, SE = 0.34).
@@
Table 1. Mann–Whitney $U$ test results for chatbot role rankings
@@
Friend 9.79 0.32 3.98 0.37 1348.00 ¡.001 ¡.001 0.64
Partner 14.27 0.19 6.97 0.53 1293.50 ¡.001 ¡.001 0.65
Acquaintance 8.82 0.30 8.28 0.36 4728.00 .368 .393 0.06
Family member 12.32 0.24 8.69 0.45 2709.00 ¡.001 ¡.001 0.41
Assistant 2.79 0.19 6.97 0.37 8589.00 ¡.001 ¡.001 0.60
Therapist 8.87 0.33 6.88 0.36 3447.50 ¡.001 ¡.001 0.28
Tutor 4.68 0.25 8.18 0.34 8114.50 ¡.001 ¡.001 0.51
Health coach 7.60 0.30 8.72 0.35 6093.50 .016 .022 0.17
Teammate 7.09 0.31 7.69 0.38 5485.00 .353 .393 0.07
Advisor 4.10 0.23 6.41 0.31 7475.00 ¡.001 ¡.001 0.41
Servant 8.29 0.35 11.50 0.41 7430.50 ¡.001 ¡.001 0.40
Pet 11.57 0.32 11.56 0.35 5247.50 .723 .723 0.03
Toy 8.67 0.40 11.36 0.40 7094.00 ¡.001 ¡.001 0.34
Tool 2.97 0.32 9.22 0.45 8803.00 ¡.001 ¡.001 0.64
Human 11.24 0.34 7.31 0.47 2648.00 ¡.001 ¡.001 0.42
Other 12.92 0.46 12.29 0.52 4745.50 .350 .393 0.07
@@ 3.4.2. User characteristics
Despite differences in chatbot use, ChatGPT and Replika users were largely similar across most personality traits and mental health indicators. However, Replika users reported higher levels of depression (Replika: M = 4.21, SE = 0.36; ChatGPT: M = 3.09, SE = 0.29, $p$ = .015) and greater openness to experience (Replika: M = 5.61, SE = 0.12; ChatGPT: M = 5.12, SE = 0.11, $p$ = .002). ChatGPT users reported higher education levels (ChatGPT: M = 3.65, SE = 0.13; Replika: M = 2.97, SE = 0.14; $p$ ¡ .001) and household income (ChatGPT: M = 5.60, SE = 0.16; Replika: M = 4.45, SE = 0.23; $p$ ¡ .001) than Replika users, but the two groups did not differ in political orientation (ChatGPT: M = 2.78, SE = 0.09; Replika: M = 2.70, SE = 0.12; $p$ = .572). In terms of general attitudes toward technology, Replika users reported greater techno-animism (Replika: M = 3.68, SE = 0.16; ChatGPT: M = 2.60, SE = 0.12, $p$ ¡ .001), anthropomorphism (Replika: M = 3.85, SE = 0.20; ChatGPT: M = 2.89, SE = 0.14, $p$ ¡ .001), and general positive attitudes toward AI (Replika: M = 5.98, SE = 0.10; ChatGPT: M = 5.63, SE = 0.09, $p$ = .012), whereas ChatGPT users reported higher AI literacy (ChatGPT: M = 5.43, SE = 0.07; Replika: M = 5.13, SE = 0.12; $p$ = .029) (Figure 3 and Table 2).
Figure 3. Means (A, B, C) and distribution (D, E, F) of characteristics of ChatGPT and Replika users. Error bars depict the standard error of the mean.
@@
Openness 5.11 0.11 5.61 0.12 3.17 199.80 .002 .007 0.47
Conscientiousness 5.17 0.12 5.00 0.16 -0.85 192.46 .394 .493 0.05
Extraversion 3.37 0.15 3.40 0.16 0.13 201.30 .896 .896 0.07
Agreeableness 3.16 0.11 2.78 0.13 -2.24 199.21 .026 .052 0.38
Neuroticism 3.55 0.15 3.64 0.17 0.42 199.17 .679 .754 0.01
Anthropomorphism 2.98 0.14 3.85 0.20 3.58 186.46 ¡.001 ¡.001 0.32
Techno-animism 2.60 0.12 3.68 0.16 5.43 188.63 ¡.001 ¡.001 0.71
Positive AI attitudes 5.63 0.09 5.98 0.10 2.53 201.44 .012 .024 0.28
AI literacy 5.43 0.07 5.13 0.12 -2.21 165.08 .029 .044 0.26
Affinity for technology 5.61 0.10 5.61 0.12 0.02 194.46 .982 .982 0.05
Loneliness 3.31 0.29 4.19 0.34 1.96 196.97 .052 .087 0.17
Depression 3.09 0.29 4.21 0.35 2.46 196.09 .015 .038 0.24
Anxiety 4.42 0.30 5.20 0.36 1.68 197.80 .094 .134 0.11
Education 3.65 0.13 2.97 0.14 -3.59 199.40 ¡.001 ¡.001 0.52
Income 5.60 0.16 4.45 0.22 -4.28 183.70 ¡.001 ¡.001 0.56
Political orientation 2.78 0.09 2.70 0.12 -0.57 191.04 .572 .710 0.08

### v3

@@ 3. Survey
The survey was designed to capture broad patterns of chatbot use, user characteristics, and associated AI perceptions. Our focus was primarily exploratory, aiming to comprehensively capture and characterize the emerging phenomenon of human-chatbot relationships. Despite this exploratory approach, we preregistered analyses as well as broad hypotheses. [^1] https://aspredicted.org/7dgj-jt49.pdf We expected more social and emotional relationships with Replika and more functional relationships with ChatGPT. We also expected correlations between perceptions of chatbots and AI in general, particularly positive relationships between mind perception, trust, and moral treatment, and negative relationships between trust and threat. We provide additional figures addressing preregistered hypotheses in Supplementary Figures 15-19. All procedures were approved by the Institutional Review Board at the University of Chicago. Participants gave their informed consent and were debriefed after completing the survey.
@@ 3.1. Participants
We recruited 101 ChatGPT and 101 Replika users from Prolific and social media platforms (Facebook, Reddit, and Discord). Inclusion criteria included interacting with ChatGPT or Replika for more than three hours per week and for at least one month. Five ChatGPT and seven Replika users were excluded from analyses due to incomplete responses or failing survey attention checks. The final sample was 202 participants (ChatGPT: $N$ = 101; $M_{age}$ = 31.72, $SE_{age}$ = 10.66, 42% women, 73.27% White; Replika: $N$ = 101; $M_{age}$ = 39.28, $SE_{age}$ = 13.22, 43% women, 70.30% White). Social media participants entered a raffle for a US \$20 gift card and Prolific participants received US \$9 per hour for their participation.
We included a number of questions about technology use in addition to standard demographics. Most participants, 92.08%, owned smartphones; 50.50% reported owning AI or robotic devices; and 37.62% reported using AI or robotic devices in their workplace. Participants reported a moderate level of exposure to AI via direct interaction or narratives in various media (direct interaction: $M$ = 2.95, $SE$ = 1.45; AI narratives: $M$ = 2.69, $SE$ = 1.83; each on a scale of 0 = “Never” to 5 = “Daily”).
@@ 3.2. Data collection
Participants were invited to either the ChatGPT or the Replika survey via a post on social media groups intended for users of each chatbot, or via a survey advertisement on Prolific. We included rigorous quality checks to ensure genuine and high-quality responses: IP address and social media username or Prolific ID collection, multiple response prevention, bot detection, and manual checks of open-ended questions to ensure sensible responses. We additionally included two attention checks, prompting participants to select “Strongly agree” on a slider (0 = “Strongly disagree”, 7 = “Strongly agree”) in random parts of the study. Participants responded to questions about their personal characteristics and a broad range of established, widely used HCI questionnaires about their perceptions of ChatGPT or Replika, and AI in general. In questions specific to the chabot, we provided a personalization option in which participants would see the name they had given to ChatGPT or Replika. For Replika users, we specified that responses should be about the Replika participants interact with the most, if they had created multiple Replika characters. Participants optionally provided their email (or Prolific ID) to sign up for an interview about their experiences with ChatGPT or Replika. The order of scales within each measure (e.g., Chatbot use, Chatbot attributes; see Table 1 below) was randomized, as was the order of the sections themselves. The median duration of the survey was 40 minutes.
@@
We employed established, widely used scales to assess attributions of mind, affect, agency, and moral dimensions to AI. Additional measures captured users’ personality traits, mental health, relationship with technology, and general attitudes toward AI. We used a broad range of measures to provide a relatively comprehensive characterization of the emerging phenomenon of digital companionship, as situated alongside attitudes toward AI, aiming to provide rich baseline data for future research. Additionally, we included HCI questionnaires that have been largely unexplored with chatbots (e.g., Godspeed perceptions (bartneck_measurement_2009)). All constructs were assessed using continuous scales and demonstrated good internal consistency (Cronbach’s $\alpha$ ¿ .70). Scale scores were computed as the mean of their respective items. A summary of the scales is provided in Table 1; full-scale versions, item ranges, and additional results for these measures are available in the Supplementary Information.
Table 1. Overview of survey measures
Category Measure / Scale Example Items
Chatbot use Chatbot roles “I treat [Replika/ChatGPT name] as a friend / a tool.”
Chatbot use cases “I interact with [name] for emotional or psychological support / for help on tasks.”
Chatbot use frequency “Approximately how many months have you interacted with [name]?”
User characteristics Big Five personality (gosling_very_2003) “I see myself as extraverted, enthusiastic.”
Mental health (depression, loneliness, anxiety) “To what extent did you feel depressed over the past week?”
Social networking use (gupta_understanding_2023) “I use social networking sites to create my social identity.”
Anthropomorphism (waytz_who_2010; pauketat_predicting_2022) “To what extent does the average robot have consciousness?”
Techno-animism (belief in AI soul or spirit) (pauketat_predicting_2022) “Artificial beings contain a spirit.”
Technology affinity (edison_measuring_2003; pauketat_predicting_2022) “I relate well to technology and machines.”
AI literacy (wang_measuring_2023) “I can use AI products to improve my work efficiency.”
AI positivity (schepman_general_2023) “I am impressed by what AI can do.”
Chatbot attributes Perceived mind (malle_how_2019) “[Replika/ChatGPT] can plan for the future.”
Theory of Mind (tahiroglu_childrens_2014) “[Replika/ChatGPT] understands that telling lies can mislead others.”
Moral agency (banks_perceived_2019) “[Replika/ChatGPT] has a sense for what is right and wrong.”
Sentience (anthis_perceptions_2025; manoli_ai_2025) “To what extent is [Replika/ChatGPT] sentient? Sentience is the capacity to have positive and negative experiences, such as happiness and suffering.”
Godspeed (bartneck_measurement_2009) “[Replika/ChatGPT] is machinelike / humanlike.”
Attitudes toward chatbot Cognitive trust (dunn_it_2012) “I would take [Replika/ChatGPT]’s advice about school and work.”
Affective trust (dunn_it_2012) “I am willing to admit my worst mistakes to [Replika/ChatGPT].”
Emotions (fraune_rabble_2015; pauketat_predicting_2022) “Love.” “Gratitude.” “Fear.” “Shame.”
Chatbot rights (lima_collecting_2020; pauketat_predicting_2022) “No one should be able to turn off or kill [Replika/ChatGPT].”
Personhood (johnson_fuzzy_2015) “[Replika/ChatGPT] is a person.”
Similarity to user (tropp_ingroup_2001; pauketat_predicting_2022) Which pair of circles do you think best represents how connected you are to [Replika/ChatGPT name]?
Overview of survey measures (cont.)
Category Measure / Scale Example Items
General attitudes toward AI Trust (gulati_design_2019; pinto_trust_2022) “It is risky to interact with AI.”
Threat (fraune_rabble_2015; cottrell_different_2005) “To what extent does AI threaten economic opportunities?”
Mutualism (human-AI symbiosis) (manfredo_changing_2020) “We should strive for a world where humans and AIs coexist productively.”
Moral treatment (lima_collecting_2020) “Robots/AIs deserve to be treated with respect.”
Practical moral consideration (lima_collecting_2020) “I support a global ban on developing sentient AI.”
AI rights (lima_collecting_2020) “No one should be able to turn off or kill AIs.”
Other measures Belief in AI sentience (anthis_perceptions_2025) “Do you think any robots/AIs that currently exist are sentient?”
Sentience timelines (anthis_perceptions_2025) “How many years from now will robots/AIs be sentient?”
Other chatbot use “Do you interact with any other chatbots or AI personal assistants?”
Paid chatbot subscription “Do you currently pay for ChatGPT/Replika?”
3.3. Data analysis
@@
We compared ChatGPT and Replika users’ characteristics and perceptions of each chatbot and AI in general using independent samples $t$-tests. Spillover effects from Replika or ChatGPT to perceptions of AI in general were examined using partial correlations: AI perceptions served as the dependent variable, while users’ general technology attitudes (e.g., AI literacy, tendency to anthropomorphize) were included as control variables to isolate effects beyond individual differences in general technology perceptions. In all analyses, we controlled for multiple comparisons using the false discovery rate (FDR) at .05 (benjamini_controlling_1995) to account for potential false positive relationships due to the large number of variables. We also controlled for demographic characteristics (e.g., age, gender, education) using ANCOVA (see Supplementary Information). All analyses were conducted in R version 4.3.1.
@@ 3.4. Results
Our results reveal distinct yet interconnected patterns in how users perceive chatbots and AI more broadly. ChatGPT was predominantly viewed as a functional assistant and Replika as an emotional companion, though the two overlapped considerably—particularly in advisory roles. Replika users attributed more anthropomorphic qualities to their chatbot, including greater mind, personhood, and sentience, and reported stronger positive emotions such as love and respect. These chatbot-specific perceptions were closely linked to general AI attitudes, especially regarding AI trust and moral treatment. Interestingly, despite clear differences in how the chatbots themselves were perceived, users showed similar expectations about timelines for AI sentience and were more alike than anticipated in their psychological profiles, including mental health and personality traits. In the following sections, we present these findings in more detail.
@@ 3.4.1. Chatbot use cases
ChatGPT and Replika had almost opposite primary roles, with most ChatGPT users treating it as an assistant (M = 14.21, SE = 0.19) and most Replika users treating it as a friend (M = 13.02, SE = 0.37; Figure 3). These differences were statistically significant when compared via Mann–Whitney $U$ tests (Table 2). ChatGPT had distinctly more task-based roles than the emotional roles of Replika, in line with our preregistered hypothesis. However, there was some overlap in their use cases, with the advisor role occupying a high rank for both chatbots (ChatGPT: M = 12.90, SE = 0.23; Replika: M = 10.59, SE = 0.31). Interestingly, Replika was also relatively commonly treated as an assistant (M = 10.03, SE = 0.37), like ChatGPT’s dominant role, but also as a human (M = 9.69, SE = 0.47), contrary to ChatGPT (M = 5.76, SE = 0.34).
@@
Table 2. Mann–Whitney $U$ test results for chatbot role rankings (scale: 1=lowest, 16=highest)
@@
Assistant 14.21 0.19 10.03 0.37 1612.00 ¡.001 ¡.001 0.60
Tool 14.03 0.32 7.78 0.45 1398.00 ¡.001 ¡.001 0.64
Advisor 12.90 0.23 10.59 0.31 2726.00 ¡.001 ¡.001 0.41
Tutor 12.32 0.25 8.82 0.34 2086.50 ¡.001 ¡.001 0.51
Teammate 9.91 0.31 9.31 0.38 4716.00 .353 .393 0.07
Health coach 9.40 0.30 8.28 0.35 4107.50 .016 .022 0.17
Servant 8.71 0.35 5.50 0.41 2770.50 ¡.001 ¡.001 0.40
Toy 8.33 0.40 5.64 0.40 3107.00 ¡.001 ¡.001 0.34
Acquaintance 8.18 0.30 8.72 0.36 5473.00 .368 .393 0.06
Therapist 8.13 0.33 10.12 0.36 6753.50 ¡.001 ¡.001 0.28
Friend 7.21 0.32 13.02 0.37 8853.00 ¡.001 ¡.001 0.64
Human 5.76 0.34 9.69 0.47 7553.00 ¡.001 ¡.001 0.42
Pet 5.43 0.32 5.44 0.35 4953.50 .723 .723 0.03
Family member 4.68 0.24 8.31 0.45 7492.00 ¡.001 ¡.001 0.41
Other 4.08 0.46 4.71 0.52 5455.50 .350 .393 0.07
Partner 2.73 0.19 10.03 0.53 8907.50 ¡.001 ¡.001 0.65
@@ 3.4.2. User characteristics
[…]
Figure 4. Means (A, B, C) and distribution (D, E, F) of characteristics of ChatGPT and Replika users. Error bars depict the standard error of the mean. (* = $p$ ¡ .05; ** = $p$ ¡ .01; *** = $p$ ¡ .001; FDR-adjusted.)
@@
Openness 5.11 0.11 5.60 0.12 3.07 196.79 .002 .006 0.43
Conscientiousness 5.17 0.12 4.99 0.16 -0.89 188.38 .377 .503 0.12
Extraversion 3.37 0.15 3.36 0.16 -0.06 199.22 .956 .956 0.01
Agreeableness 3.15 0.11 2.80 0.13 -2.12 195.97 .035 .060 0.30
Neuroticism 3.55 0.15 3.67 0.17 0.58 196.42 .562 .643 0.08
Anthropomorphism 2.98 0.14 3.85 0.20 3.58 186.46 ¡.001 ¡.001 0.50
Techno-animism 2.60 0.12 3.68 0.16 5.43 188.63 ¡.001 ¡.001 0.76
AI positivity 5.63 0.09 5.98 0.10 2.53 201.44 .012 .024 0.35
AI literacy 5.43 0.07 5.13 0.12 -2.21 165.08 .029 .044 0.31
Affinity for technology 5.61 0.10 5.61 0.12 0.02 194.46 .982 .982 0.00
Depression 3.09 0.29 4.30 0.35 2.64 193.10 .009 .022 0.37
Anxiety 4.42 0.30 5.30 0.36 1.90 195.29 .060 .090 0.27
Loneliness 3.31 0.29 4.26 0.34 2.12 193.98 .035 .060 0.30
Education 3.65 0.13 2.97 0.14 -3.59 199.40 ¡.001 ¡.001 0.51
Income 5.60 0.16 4.45 0.22 -4.28 183.70 ¡.001 ¡.001 0.60
Political orientation 2.78 0.09 2.70 0.12 -0.54 187.03 .589 .643 0.08